# Right on Time: Revising Time Series Models by Constraining their Explanations

### v2

@@ 4.1 Experimental Setup
Data. We perform experiments on various different datasets. For classification, we focus mainly on the UCR/UEA repository (Dau et al., 2018), which holds a wide variety of datasets for this task. The data originates from different domains, e.g., health records, industrial sensor data, and audio signals. We select all available datasets of a minimal size (cf. App. A.2), which results in Fault Detection A, Ford A, Ford B, and Sleep. We omit experiments on the very small datasets of UCR, as these are generally less suited for deep learning (Ismail Fawaz et al., 2020). For time series forecasting, we evaluate on three popular datasets of the Darts repository (Herzen et al., 2022): ETTM1, Energy and Weather, which all have been used with the TiDE model before (Das et al., 2023). These datasets are sufficiently large, allowing us to investigate the impact of confounding behavior in isolation without the risk of overfitting. We standardize all datasets as suggested by Wu et al. (2021), i.e., rescaling the distribution of values to zero mean and a standard deviation of one.
Production Press Sensor Data (P2S). Our method aims to mitigate real-world confounders in time series data. For a more accurate and reliable assessment of the method’s performance, it is crucial to consider the variations and noise of real-world data and, importantly, real-world confounders. However, there are no available datasets showcasing this problem, challenging models to overcome their confounders. Thus, we introduce the dataset Production Press Sensor Data (P2S) to fill this gap. [^2] https://huggingface.co/datasets/AIML-TUDA/P2S
P2S consists of sensor recordings from a production line for sheet metal parts using a high-speed press. This procedure is one of the sheet metal working industry’s most economically significant manufacturing processes. Specifically, we consider the sensor recordings of the deep bowing phase. Based on the sensor data, the task is to predict a run is normal or has a defect. Experiment runs are conducted with different production speeds, which does not affect the occurrence of defects. However, the speed affects the process’s friction and, thus, the system’s applied forces, leading to confounding factors in the sensor data.
In Fig. 3, this can be clearly seen when comparing the sensor recordings for normal and defect runs at speeds of 80 or 225 strokes per minute. The actual differences between normal and defect runs at the same speed are more subtle; for example, the minimal force can be an indicator. The challenge of the dataset for a model is to pick up on these factors, instead of the deflecting characteristics of the stroke rate. In particular, if no faulty and normal runs of the same speed are in the training data, there is a strong potential for models to focus only on the confounding factors and thus fail to generalize beyond training. Investigating how models can or cannot avoid picking up speed as a confounding factor provides valuable insights into how robust they are to such challenges in the real world. With P2S, we provide a first real-world data basis to evaluate models for their sensitivity to confounding factors in time series data.
@@
Figure 3: Samples of P2S with normal (top) and defect (bottom) setting at 80 and 225 strokes per minute. Areas that vary depending on the stroke rate and are considered confounding and marked red.
The dataset consists of 2264 samples, divided into roughly 200 normal and defect runs of six different speeds. Each of the samples has sensor data of 4096 time steps. In the default setting, the training data has normal and defect runs of different speeds. A reference setting where the model is not affected by any confounder is curated by having normal and defect runs of specifically the same speed in the training set. Further details on the dataset can be found in App. A.4.
Models. For the time series classification datasets, we use the FCN model of Ma et al. (2023). On the dataset Sleep, which was added to UCR after their study, we slightly modified their architecture, which achieves better results (cf. A.1). For forecasting, we use the recently introduced TiDE model (Das et al., 2023). Both models achieve SOTA or at least comparable performance on the respective datasets.
@@
Metrics. In our evaluations, we compare the performance of models on confounded and unconfounded datasets with and without RioT. For classification, we use balanced (multiclass) accuracy (ACC), and for forecasting, we use mean squared error (MSE) and mean absolute error (MAE).
@@
Table 1: Applying RioT mitigates confounders in time series classification. Performance of the FCN model before and after applying RioT for spatial (FCNsp) and frequency (FCNfreq) confounders separately. FCNReference represents the ideal scenario where the model is not affected by any confounder. The high train accuracy and low test accuracy for FCN shows overfitting to the confounder. Applying RioT improves the test performance, indicating its effectiveness in reducing the influence of individual confounders (mean and std over 5 runs).
@@
Config Fault Detection Ford A Ford B Sleep
@@
FCNReference 0.99 ±0.00 0.99 ±0.00 0.92 ±0.01 0.91 ±0.00 0.93 ±0.00 0.76 ±0.01 0.68 ±0.00 0.62 ±0.00
FCNsp 1.00 ±0.00 0.74 ±0.06 1.00 ±0.00 0.71 ±0.08 1.00 ±0.00 0.63 ±0.03 1.00 ±0.00 0.10 ±0.03
FCNsp + RioTsp 0.99 ±0.01 0.92 ±0.03 0.99 ±0.01 0.84 ±0.02 0.99 ±0.00 0.68 ±0.02 0.60 ±0.06 0.54 ±0.05
FCNfreq 0.98 ±0.01 0.87 ±0.03 0.98 ±0.00 0.73 ±0.01 0.99 ±0.01 0.60 ±0.01 0.98 ±0.00 0.27 ±0.02
FCNfreq + RioTfreq 0.94 ±0.00 0.90 ±0.03 0.91 ±0.05 0.77 ±0.00 0.97 ±0.00 0.62 ±0.00 0.67 ±0.05 0.44 ±0.06
Table 2: RioT can successfully overcome confounders in time series forecasting. MAE/MSE values on the confounded training set and the unconfounded test set with TiDEReference being the ideal scenario where the model is not affected by any confounder. TiDE shows degraded performance on spatial and frequency confounded datasets, which are mitigated by applying RioT (mean and std over 5 runs).
Metric Config ETTM1 Energy Weather
@@
$MAE\downarrow$ TiDEReference 0.36 ±0.01 0.48 ±0.01 0.40 ±0.01 0.38 ±0.02 0.16 ±0.01 0.11 ±0.02
TiDEsp 0.32 ±0.01 0.54 ±0.01 0.40 ±0.01 0.85 ±0.01 0.18 ±0.01 0.28 ±0.02
TiDEsp + RioTsp 0.34 ±0.01 0.51 ±0.01 0.57 ±0.01 0.58 ±0.01 0.22 ±0.01 0.18 ±0.00
TiDEfreq 0.18 ±0.01 0.74 ±0.06 0.18 ±0.01 0.53 ±0.07 0.20 ±0.01 0.36 ±0.06
TiDEfreq + RioTfreq 0.19 ±0.01 0.60 ±0.05 0.18 ±0.01 0.40 ±0.03 0.20 ±0.01 0.35 ±0.02
$MSE\downarrow$ TiDEReference 0.27 ±0.01 0.47 ±0.01 0.27 ±0.01 0.26 ±0.02 0.05 ±0.01 0.03 ±0.01
TiDEsp 0.22 ±0.01 0.54 ±0.03 0.28 ±0.01 1.19 ±0.03 0.07 ±0.01 0.14 ±0.02
TiDEsp + RioTsp 0.23 ±0.01 0.48 ±0.01 0.53 ±0.02 0.52 ±0.02 0.10 ±0.01 0.06 ±0.00
TiDEfreq 0.06 ±0.01 0.69 ±0.08 0.07 ±0.01 0.34 ±0.08 0.09 ±0.01 0.16 ±0.04
TiDEfreq + RioTfreq 0.07 ±0.01 0.49 ±0.07 0.07 ±0.01 0.21 ±0.02 0.09 ±0.01 0.15 ±0.02
Confounders. To evaluate how well RioT can mitigate confounders in a more controlled setting, we add shortcuts to the datasets from the UCR and Darts repositories. For classification, we add spatial (FCNsp) or frequency (FCNfreq) shortcut confounders to each sample. The spatial confounder is a distinct pattern per class, replacing the original time series at affected time points. The frequency confounder is a unique frequency per class, which is added to the data, affecting all steps of the time series. The spatial confounder for forecasting (TiDEsp) is also a shortcut confounder, where we copy the forecasting window back into the lookback. This results in a shortcut for every second sample and minor perturbations for the remaining data points. The frequency confounder (TiDEfreq) consists of reoccurring noise in the form of Dirac impulses added to the data. This is not a shortcut confounder but simply distracts from the forecasting task. Confounders are only in the training data, not validation or test data. We generate an annotation mask based on the confounder area or frequency to simulate human feedback. This mask is applied to all confounded samples.
4.2 Evaluations
Removing Confounders for Time Series Classification. First, we evaluate the effectiveness of RioT (RioTsp with spatial and RioTfreq with frequency feedback) in revising confounders in classification tasks. We compare the balanced (multiclass) accuracy of the FCN model in two scenarios: without the application of RioT and with it. As shown in Tab. 1, (top), FCN without RioT overfits to the shortcuts, achieving 100% training accuracy on all datasets but fails to maintain this on the test set. In contrast, applying RioT considerably improves the model’s test performance across all datasets. Although it does not fully reach the upper bound performance of the reference scenario where FCN is trained on unconfounded data, RioT improves performance on average by more than 10%. Additionally, the unconfounded FCN has a notable drop in training-to-test performance on the Ford B dataset, suggesting a shift between training and test distributions. Nevertheless, using RioT enhances the model’s performance on this dataset, too.
Further evaluations of FCN, both with and without RioTfreq, on frequency-confounded data (Tab. 1, bottom) exhibit a pattern similar to that seen with spatial confounders. The model learns the training data shortcuts, achieving near-perfect accuracy, which does not translate to the test set. RioTfreq with frequency feedback again elevates the test performance for all datasets. The least pronounced improvement is on Ford B, with only about a 1% increase when using RioTfreq, leading to the hypothesis that the confounded frequencies in this dataset may carry crucial information, which is inadvertently obscured by the application of RioTfreq.
Removing Confounders for Time Series Forecasting. Confounders are not exclusive to time series classification and can significantly impact other tasks, such as forecasting. In Tab. 2, we demonstrate that confounding factors in the data drastically affect model generalization. The spatially confounded model (TiDEsp) exhibits a notable discrepancy between training and test performance, suggesting it overfitts to the spatial shortcuts. Applying RioTsp to the model lowers both test MAE and MSE on all datasets while having a particularly pronounced effect for Energy of approximately 56% lower test MSE.
In the frequency-confounded setting, the training data includes a recurring noise as a distracting confounder. This again lowers the test performance of TiDE substantially as the model overfits to the training pattern and then fails to generalize. In contrast, using RioTfreq mitigates the model’s excessive focus on the frequency confounder, improving test performances. For instance, in the ETTM1 dataset, the use of RioTfreq decreases the test MAE from 0.54 to 0.51 (a reduction of 5.5%). Similarly, the test MSE decreases from 0.54 to 0.48, representing an 11.1% decrease.
These outcomes from spatial and frequency confounders highlight the efficacy of RioT in addressing different types of confounding factors in time series forecasting. Overall, applying RioT substantially improves test performance for both types of confounders. In some cases, it even reaches the results of the unconfounded reference model, highlighting that RioT fully overcomes the confounder. Interestingly, on the Energy dataset, the performance with RioTfreq even surpasses the reference model. Here, the added frequency has the effect of a form of data augmentation, enhancing the model’s robustness.
@@ 4.2 Evaluations
Table 3: Applying RioT overcomes the confounder in P2S. Accuracy values on the confounded train set and the unconfounded test set. FCN learns the train confounder, resulting in a drop in test performance. Applying RioT with partial feedback (2) already yields good improvements, while adding feedback on the full confounder area (4) is even better. FCNReference is the ideal scenario where the data is specifically curated so that FCN is not affected by any confounder (mean and std over 5 runs).
P2S Train Test
FCNReference 0.97 ±0.01 0.95 ±0.01
FCN 0.99 ±0.01 0.66 ±0.14
FCN + RioTsp (2) 0.96 ±0.01 0.78 ±0.05
FCN + RioTsp (4) 0.95 ±0.01 0.82 ±0.06
@@
Table 4: RioT can combine spatial and frequency feedback. If the data is confounded in the time and frequency domain, RioT can combine feedback on both domains to mitigate confounders. This is superior to applying feedback on only one domain. Reference values represent the ideal scenario when the model is not affected by any confounder (mean and std over 5 runs).
@@
TiDEfreq,sp + RioTfreq,sp 0.25 ±0.01 0.47 ±0.01
Removing Confounders in the Real-World. So far, our experiments have demonstrated the ability to counteract confounders within controlled environments where the confounders were intentionally embedded in the data. However, real-world scenarios often lack this known confounder structure. Our new proposed dataset P2S presents a scenario under real-world conditions. The model explanations for a sample in Fig. 4 (top) reveal a primary focus on four distinct regions of the sensor curve. Using this explanation, one can identify that the model focuses on two regions in the middle. With domain knowledge, it is then possible to state that these regions should not be relevant for predicting whether the outcome is normal or defective. By applying RioT, we can redirect the model’s attention away from these regions. Tab. 3 shows that the initial model is indeed focusing on the confounding run speed, resulting in high training and substantially lower test accuracy. Using RioT to apply feedback to two initially discovered areas already improves test performance substantially (2). However, the explanations of Fig. 4 (middle) reveal that the model still focuses on some confounded input parts. Applying feedback on all four detected regions with RioT improves the test performance even more Tab. 3 (4). While it does not reach the reference setting, where the dataset was specifically curated so that the model does not pick up on the speed confounder in training, applying RioT with two and all four feedback regions shows considerable improvements. Additionally, these results highlight exemplarily the effectiveness of RioT in real-world scenarios where potentially not the full confounder is known (from the beginning).
On P2S, the model has a large variation in test performance, as indicated by the standard deviation in Tab. 3. The variation persists, albeit reduced, even after applying RioT. This suggests intrinsic complexity and variability of how the model learns and adapts to confounders. Nonetheless, applying feedback via RioT consistently narrows the standard deviation, enhancing model reliability and decision-making quality in real-world applications.
Removing Multiple Confounders at Once. In the previous experiments, we illustrated that RioT is suitable for addressing individual confounding factors, whether spatial or frequency-based. Real-world time series data, however, often present a blend of multiple confounding factors that simultaneously may influence model performance. Therefore, we explore the effects of applying RioT simultaneously to spatial and frequency confounders in Tab. 4. When Sleep is confounded both in the spatial and frequency domain, FCN without RioT overfits to the confounders and does not generalize to the test set. Applying only spatial or frequency feedback does not mitigate the confounders, as the model simply learns the other confounder. Combining feedback in both domains (RioTfreq,sp), on the other hand, shows good improvements on the test set, on par with the performance of the previously identified frequency-confounded scenario (cf. Tab. 1). This result aligns with our expectations; given that the original dataset was confounded, the model’s ability to utilize information was limited by the confounders we introduced. These insights not only confirm the utility of RioTfreq,sp in multi-confounder contexts but also suggest a performance upper limit for performance when the underlying data is inherently confounded.
Tab. 4 (bottom) illustrates the impact of multiple confounders on the Energy dataset in the context of forecasting. It is clearly visible that when faced with both spatial shortcut and noise confounders, the model tends to overfit the training data, as indicated by the lower training MSE (corresponding MAE is detailed in the Appendix in Tab. 5). Applying either spatial or frequency feedback individually results in some improvement, as either the spatial shortcut or the frequency noise is avoided. Applying both types of feedback simultaneously (RioTfreq,sp) demonstrates the largest improvements on the test set as feedback on both confounder types is present at the same time.
Nonetheless, the distance between the performance of RioTfreq,sp and the non-confounded model is more pronounced than in the single confounder cases (cf. Tab. 2), suggesting a compounded challenge. This gap indicates the complex nature of interacting confounders, which appears to create a performance upper limit in this case. While RioT is suitable for reducing the confounding effects, the tangled interplay between spatial and frequency confounders may limit the extent to which we can replicate the ideal unconfounded scenario. Research to optimize the deconfounding process in more complex data environments thus represents an important future step.

### v3

@@ 4.1 Experimental Setup
Data. We perform experiments on various datasets. For classification, we focus mainly on the UCR/UEA repository (Dau et al., 2018), which holds a wide variety of datasets for this task. The data originates from different domains, e.g., health records, industrial sensor data, and audio signals. We select all available datasets of a minimal size (cf. SubSec. A.3), which results in Fault Detection A, Ford A, Ford B, and Sleep. We omit experiments on the very small datasets of UCR, as these are generally less suited for deep learning (Ismail Fawaz et al., 2020). We use the splits provided by the UCR archive. For time series forecasting, we evaluate on three popular datasets of the Darts repository (Herzen et al., 2022): ETTM1, Energy, and Weather with 70%/30% train/test splits. These datasets are sufficiently large, allowing us to investigate the impact of confounding behavior in isolation without the risk of overfitting. We standardize all datasets as suggested by Wu et al. (2021), i.e., rescaling the distribution of values to zero mean and a standard deviation of one.
Production Press Sensor Data (P2S). RioT aims to mitigate confounders in time series data. To assess our method, we need datasets with annotated real-world confounders. So far, there are no such datasets available. To fill this gap, we introduce Production Press Sensor Data (P2S) [^2] https://huggingface.co/datasets/AIML-TUDA/P2S, a dataset of sensor recordings with naturally occurring confounders. The sensor data comes from a high-speed press production line for metal parts, one of the sheet metal working industry’s most economically significant processes. The task is to predict whether a run is defective based on the sensor data. The recordings include different production speeds, which, although not affecting part quality, influence process friction and applied forces. Fig. 3 shows samples recorded at different speeds from normal and defect runs, highlighting variations even within the same class. An expert identified regions in the time series that vary with production speed, potentially distracting models from relevant classification indicators, especially when no defect and normal runs of the same speed are in the training data. Thus, the run’s speed is a confounder, challenging models to generalize beyond training. The default P2S setting includes normal and defect runs of different speeds, with an unconfounded setting of runs at the same speed. Further details on the dataset are available in App. B.
@@
Figure 3: Samples of P2S with normal (left) and defect (right) setting at 80 and 225 strokes per minute. Areas that vary depending on the stroke rate and are considered confounding and marked red.
Models. For time series classification, we use the FCN model of Ma et al. (2023), with a slightly modified architecture for Sleep to achieve a better unconfounded performance (cf.  SubSec. A.2). Additionally, we use the OFA model by Zhou et al. (2023). For forecasting, we use the recently introduced TiDE model (Das et al., 2023), PatchTST (Nie et al., 2023) and NBEATS (Oreshkin et al., 2020) to highlight the applicability of our method to a variety of model classes.
@@
Metrics. In our evaluations, we compare the performance of models on confounded and unconfounded datasets with and without RioT. For classification, we report balanced (multiclass) accuracy (ACC), and for forecasting the mean squared error (MSE). The corresponding mean absolute error (MAE) results can be found in SubSec. A.6. We report average and standard deviation over 5 runs.
@@
Table 1: Applying RioT mitigates confounders in time series classification. Performance before and after applying RioT for spatial (SP Conf) and frequency (Freq Conf) confounders separately. Unconfounded represents the ideal scenario where the model is not affected by any confounder.
@@
Model Config (ACC $\uparrow$) Fault Detection A FordA FordB Sleep
@@
FCN Unconfounded 0.99 ±0.00 0.99 ±0.00 0.92 ±0.01 0.91 ±0.00 0.93 ±0.00 0.76 ±0.01 0.68 ±0.00 0.62 ±0.00
SP Conf 1.00 ±0.00 0.74 ±0.06 1.00 ±0.00 0.71 ±0.08 1.00 ±0.00 0.63 ±0.03 1.00 ±0.00 0.10 ±0.03
+ RioTsp 0.98 ±0.01 $\bullet$ 0.93 ±0.03 0.99 ±0.01 $\bullet$ 0.84 ±0.02 0.99 ±0.00 $\bullet$ 0.68 ±0.02 0.60 ±0.06 $\bullet$ 0.54 ±0.05
Freq Conf 0.98 ±0.01 0.87 ±0.03 0.98 ±0.00 0.73 ±0.01 0.99 ±0.01 0.60 ±0.01 0.98 ±0.00 0.27 ±0.02
+ RioTfreq 0.94 ±0.00 $\bullet$ 0.90 ±0.03 0.83 ±0.02 $\bullet$ 0.83 ±0.02 0.94 ±0.00 $\bullet$ 0.65 ±0.01 0.67 ±0.05 $\bullet$ 0.45 ±0.07
OFA Unconfounded 1.00 ±0.00 0.98 ±0.02 0.92 ±0.01 0.87 ±0.04 0.95 ±0.01 0.70 ±0.04 0.69 ±0.00 0.64 ±0.01
SP Conf 1.00 ±0.00 0.53 ±0.02 1.00 ±0.00 0.50 ±0.00 1.00 ±0.00 0.52 ±0.01 1.00 ±0.00 0.21 ±0.05
+ RioTsp 0.96 ±0.08 $\bullet$ 0.98 ±0.01 0.92 ±0.03 $\bullet$ 0.85 ±0.02 0.94 ±0.01 $\bullet$ 0.65 ±0.04 0.52 ±0.22 $\bullet$ 0.58 ±0.05
Freq Conf 1.00 ±0.00 0.72 ±0.02 1.00 ±0.00 0.65 ±0.01 1.00 ±0.00 0.56 ±0.02 0.99 ±0.00 0.24 ±0.03
+ RioTfreq 0.96 ±0.02 $\bullet$ 0.98 ±0.02 0.78 ±0.04 $\bullet$ 0.85 ±0.04 1.00 ±0.00 $\bullet$ 0.64 ±0.03 0.50 ±0.16 $\bullet$ 0.49 ±0.04
Confounders. To evaluate how well RioT can mitigate confounders in a more controlled setting, we add spatial (sp) or frequency (freq) shortcuts to the datasets from the UCR and Darts repositories. These confounders create spurious correlations between patterns and class labels or forecasting signals in the training data, but are absent in validation or test data. We generate an annotation mask based on the confounder area or frequency to simulate human feedback. More details on the confounders can be found in SubSec. A.5.
4.2 Evaluations
Removing Confounders for Time Series Classification. We evaluate the effectiveness of RioT (spatial: RioTsp, frequency: RioTfreq) in addressing confounders in classification tasks by comparing balanced accuracy with and without RioT. As shown in Tab. 1, without RioT, both FCN and OFA overfit to shortcuts, achieving $\approx$100training accuracy, while having poor test performance. Applying RioT significantly improves test performance for both models across all datasets. In some cases, RioT even reaches the performance of the ideal reference (unconfounded) scenario as if there would be no confounder in the data. Even on FordB, where the drop in training-to-test performance of the reference indicates a distribution shift, RioTsp is still beneficial. Similarly, RioTfreq enhances performance on frequency-confounded data, though the improvement is less pronounced for FCN on Ford B, suggesting essential frequency information is sometimes obscured by RioTfreq. In summary, RioT (both RioTsp and RioTfreq) successfully mitigates confounders, enhancing test generalization for FCN and OFA models.
Removing Confounders for Time Series Forecasting. Confounders are not exclusive to time series classification and can significantly impact other tasks, such as forecasting. In Tab. 2 we outline that spatial confounders cause models to overfit, but applying RioTsp reduces MSE across datasets, especially for Energy, where MSE drops by up to 56%. In the frequency-confounded setting, the training data includes a recurring Dirac impulse as a distracting confounder (cf. SubSec. A.5 for details). RioTfreq alleviates this distraction and improves the test performance significantly. For example, TiDE’s test MSE on ETTM1 decreases by 14% compared to the confounded model.
In general, RioT effectively addresses spatial and frequency confounders in forecasting tasks. Interestingly, for TiDE on the Energy dataset, the performance with RioTfreq even surpasses the unconfounded model. Here, the added frequency acts as a form of data augmentation, enhancing model robustness. A similar behavior can also be observed for NBEATS and ETTM1, where the confounded setting actually improves the model slightly, and RioT even improves upon that.
Table 2: RioT can successfully overcome confounders in time series forecasting. MSE values (MAE values cf. Tab. 7) on the confounded training set and the unconfounded test set with Unconfounded being the ideal scenario where the model is not affected by any confounder.
Model Config (MSE $\downarrow$) ETTM1 Energy Weather
@@ 4.2 Evaluations
NBEATS Unconfounded 0.30 ±0.02 0.47 ±0.02 0.34 ±0.03 0.26 ±0.02 0.08 ±0.01 0.03 ±0.01
SP Conf 0.24 ±0.01 0.55 ±0.01 0.33 ±0.03 0.94 ±0.02 0.09 ±0.01 0.16 ±0.04
+ RioTsp 0.30 ±0.01 $\bullet$ 0.50 ±0.01 0.45 ±0.03 $\bullet$ 0.58 ±0.01 0.11 ±0.01 $\bullet$ 0.09 ±0.02
Freq Conf 0.30 ±0.02 0.46 ±0.01 0.33 ±0.04 0.36 ±0.04 0.11 ±0.02 0.32 ±0.09
+ RioTfreq 0.31 ±0.02 $\bullet$ 0.45 ±0.01 0.33 ±0.04 $\bullet$ 0.34 ±0.04 0.81 ±0.48 $\bullet$ 0.17 ±0.01
PatchTST Unconfounded 0.46 ±0.03 0.47 ±0.01 0.26 ±0.01 0.23 ±0.00 0.26 ±0.03 0.08 ±0.01
SP Conf 0.40 ±0.02 0.55 ±0.01 0.29 ±0.01 0.96 ±0.03 0.20 ±0.03 0.19 ±0.01
+ RioTsp 0.40 ±0.03 $\bullet$ 0.53 ±0.01 0.44 ±0.00 $\bullet$ 0.45 ±0.01 0.55 ±0.20 $\bullet$ 0.14 ±0.01
Freq Conf 0.45 ±0.03 0.91 ±0.16 0.04 ±0.00 0.53 ±0.05 0.63 ±0.09 0.24 ±0.04
+ RioTfreq 0.91 ±0.07 $\bullet$ 0.66 ±0.04 2.45 ±4.59 $\bullet$ 0.38 ±0.06 0.96 ±0.02 $\bullet$ 0.17 ±0.00
TiDE Unconfounded 0.27 ±0.01 0.47 ±0.01 0.27 ±0.01 0.26 ±0.02 0.25 ±0.02 0.03 ±0.00
SP Conf 0.22 ±0.01 0.54 ±0.03 0.28 ±0.01 1.19 ±0.03 0.22 ±0.03 0.15 ±0.01
+ RioTsp 0.23 ±0.01 $\bullet$ 0.48 ±0.01 0.53 ±0.02 $\bullet$ 0.52 ±0.02 0.25 ±0.03 $\bullet$ 0.11 ±0.01
Freq Conf 0.06 ±0.01 0.69 ±0.08 0.07 ±0.01 0.34 ±0.08 0.79 ±0.09 0.31 ±0.09
+ RioTfreq 0.07 ±0.01 $\bullet$ 0.49 ±0.07 0.07 ±0.01 $\bullet$ 0.21 ±0.02 1.12 ±0.36 $\bullet$ 0.22 ±0.01
Removing Confounders in the Real-World. So far, our experiments have demonstrated the ability to counteract confounders within controlled environments. However, real-world scenarios often have more complex confounder structures. Our new proposed dataset P2S presents such real-world conditions. The model explanations for a sample in Fig. 4 (top) reveal a focus on distinct regions of the sensor curve, specifically the two middle regions. With domain knowledge, it’s clear that these regions shouldn’t affect the model’s output. By applying RioT, we can redirect the model’s attention away from these regions. New model explanations highlight that the model still focuses on incorrect regions, which can be mitigated by extending the annotated area. In Tab. 4, the model performance (exemplarly with FCN) in these settings is presented. Without RioT, the model overfits to the confounder. the test performance improves already with partial feedback (2) and improves even more with full feedback (4). These results highlight the effectiveness of RioT in real-world scenarios, where not all confounders are initially known.
Removing Multiple Confounders at Once. In the previous experiments, we illustrated that RioT is suitable for addressing individual confounding factors, whether spatial or frequency-based. Real-world time series data, however, often present a blend of multiple confounding factors that simultaneously may influence model performance.
@@
Table 3: Applying RioT overcomes the confounder in P2S. Performance on confounded train set and the unconfounded test set. FCN learns the train confounder, resulting in a drop in test performance. Applying RioT with partial feedback (2) already yields good improvements, while adding feedback on the full confounder area (4) is even better. Unconfounded is the ideal scenario, specifically curated so that there is no confounder.
P2S (ACC $\uparrow$) Train    Test
FCNUnconfounded 0.97 ±0.01 0.95 ±0.01
FCNsp 0.99 ±0.01 0.66 ±0.14
FCNsp + RioTsp (2) 0.96 ±0.01 0.78 ±0.05
FCNsp + RioTsp (4) 0.95 ±0.01 $\bullet$ 0.82 ±0.06
@@
Table 4: RioT can combine spatial and frequency feedback. If the data is confounded in the time and frequency domain, RioT can combine feedback on both domains to mitigate confounders, superior to feedback on only one domain. Unconfounded represents the ideal scenario when the model is not affected by any confounder.
@@
TiDEfreq,sp + RioTfreq,sp 0.25 ±0.01 $\bullet$ 0.47 ±0.01
We thus investigate the impact of applying RioT to both spatial and frequency confounders simultaneously (cf. Tab. 4), exemplary using FCN and TiDE. When Sleep is confounded in both domains, FCN without RioT overfits and fails to generalize. Addressing only one confounder does not mitigate the effects, as the model adapts to the other. However, combining feedback for both domains (RioTfreq,sp) significantly improves test performance, matching the frequency-confounded scenario (cf. Tab. 1). Tab. 4 (bottom) shows the impact of multiple confounders on the Energy dataset for forecasting. When faced with both spatial shortcut and noise confounders, the model overfits, indicated by lower training MSE. While applying either spatial or frequency feedback individually already shows some effect, utilizing both types of feedback simultaneously (RioTfreq,sp) results in the largest improvements, as both confounders are addressed. The performance gap between RioTfreq,sp and the non-confounded model is more pronounced than in single confounder cases (cf. Tab. 2), suggesting a compounded challenge. Optimize the deconfounding process in highly complex data environments thus remains an important challenge.
Figure 5: RioT uses feedback efficiently. Even with feedback on only a small percentage of the data, RioT can overcome confounders.
Feedback Generalization. As human feedback is an essential aspect of RioT, we investigate the required annotations and the potential to generalize annotations across samples. Our findings indicate that not every sample needs annotation. Fig. 5 shows that we can significantly reduce the amount of annotated data for classification and forecasting (cf. App. Tab. 6 and Tab. 5 for results on the other datasets). Even minimal feedback, such as annotating just 5% of the samples, substantially improves performance compared to no feedback. Furthermore, the results on P2S highlights that annotations can be generalized across multiple samples. Once the confounder on P2S has been identified on a couple of samples, the expert annotations can be used on full dataset. The systematic nature of shortcut confounders suggest that generalizing annotations is an effective possibility to obtain feedback efficiently. While RioT does rely on human annotations, these findings highlight that it can work without extensive manual human interactions, and that obtained annotations can be utilized efficiently.
Limitations. An important aspect of RioT is the human feedback provided in the Obtain step. Integrating human feedback into the model is a key advantage of RioT, but can also be a limitation. While we have shown that a small fraction of samples with annotations can be sufficient, and that annotations can be applied for many samples, they are still necessary for RioT. Additionally, like many other (explanatory) interactive learning methods, RioT assumes correct human feedback. Thus, considering possible repercussions of inaccurate feedback when applying RioT in practice is important. Another potential drawback of RioT are increased training costs. RioT requires computation of a mixed-partial derivative to optimize the model’s explanation when using gradient-based attributions. While this affects training cost, the loss can be formulated as a Hessian-vector product, which is fast to compute in practice, making the additional overhead easy to manage.